Find the slope of a line perpendicular to the line passing through the given two points. (- 7,6) and (0,4)
Answer by mananth(16946)

(- 7,6) and (0,4)
(x1, y1) and (x2,y2)
m= (y2-y1)/(x2-x1)
The slope of the line is (4-6)/(0-(-7))
=-2/7
the perpendicular line will have a slope = 7/2
m1*m2=-1